OpenStudy (aaronq):

\(6^x+6^{-x}=2\) taking ln doesn't work, \(xln(6)-xln(6)=ln(2)\) i end up dividing by zero. Can i do this? \(6^{|x|}+6^{|-x|}=|2|\) \(6^x+6^x=2\rightarrow 2*6^x=2\) \(6^x=1\) \(xln(6)=ln(1)\) \(xln(6)=0 \rightarrow x=\dfrac{0}{ln(6)}=0\) or am i breaking rules?

OpenStudy (lukebluefive):

I don't think you can arbitrarily take the absolute value of both sides of an equation. Otherwise: \[x = -3\] Would become: \[\left| x \right| = \left| -3 \right|\] Which results in: \[x = 3\]

OpenStudy (aaronq):

yeah, i wasn't sure about that. Do you know how to solve this otherwise? I'm probably overlooking something.

OpenStudy (lukebluefive):

My first guess was that ln should be the first thing to try, and you definitely got a valid answer. I think there's a power rule that could help here; let me find it.

OpenStudy (aaronq):

yeah, i tried taking ln, but it seems to be a dead end because i end up with \(ln\dfrac{6}{6}\) in the denominator. Alright, thanks.

OpenStudy (lukebluefive):

What about: \[a^k = b\] \[a = b^\frac{1}{k}\]

OpenStudy (lukebluefive):

Oh! \[6^x +6^{-x} = 2\] Equals: \[6^x + \frac{1}{6^x} = 2\]

OpenStudy (lukebluefive):

Rewriting: \[\frac{6^x}{1} + \frac{1}{6^x} = 2\]

OpenStudy (lukebluefive):

Multiply top and bottom by a clever form of 1: \[\frac{6^x}{1}*\frac{6^x}{6^x} + \frac{1}{6^x} = 2\]

OpenStudy (aaronq):

do you then end up with 2xln(6)=2xln(6) 2x=2x(ln6/ln6)=0

OpenStudy (lukebluefive):

Oops: \[\frac{6^{2x}}{6^x} + \frac{1}{6^x} = 2\]

OpenStudy (lukebluefive):

I'm not sure... where'd you get the other 2x?

OpenStudy (aaronq):

when i brought the \(6^x\) over, the 2 was already there

OpenStudy (aaronq):

so \(6^{2x}+1=2*6^x\)

OpenStudy (lukebluefive):

Adding: \[\frac{6^{2x} + 1}{6^x} = 2\]

OpenStudy (lukebluefive):

Lovely, okay, what did you do next?

OpenStudy (aaronq):

took ln \(2xln(6)+ln(1)=2xln(6)\) \(ln(1)=0\) wait i just noticed i end up with \(RS=2x\dfrac{ln(6)}{ln(6)}=2x\) and not zero.

OpenStudy (lukebluefive):

Is that good or bad? I'm sorry, you lost me at the RS.

OpenStudy (aaronq):

sorry, so \(2xln(6)+ln(1)=2xln(6)\) \(2xln(6)=2xln(6)\) i don't see how i can get a value for x from here, if i isolate the x's then i end losing them. \(\dfrac{x}{x}=\dfrac{ln(6)}{ln(6)}\)

OpenStudy (lukebluefive):

Did we make a mistake somewhere? Oh, I think I found it!

OpenStudy (lukebluefive):

\[\ln(xy) = \ln(x) + \ln(y)\]

OpenStudy (aaronq):

for this part? \(ln(6)+ln(1)=ln(6*1)\)

OpenStudy (aaronq):

we'll thanks for you help i appreciate it!

OpenStudy (aaronq):

i'll figure it out later

OpenStudy (lukebluefive):

Yeah, that part. Sorry, internet d/ced.

OpenStudy (lukebluefive):

So, continuing from just before we took the natural log: \[\ln(6^{2x}) + \ln(1) = \ln(2*6^x)\]

OpenStudy (lukebluefive):

This equals: \[2x*\ln(6) + 0 = \ln(2) + \ln(6^x)\]

OpenStudy (lukebluefive):

Simplifying: \[2x * \ln(6) = \ln(2) + x*\ln(6)\]

OpenStudy (lukebluefive):

I'm not sure what that gets us, though.

OpenStudy (lukebluefive):

Oh, subtract x * ln(6) from both sides: \[2x*\ln(6) - x*\ln(6) = \ln(2)\] Simplify: \[x*\ln(6) = \ln(2)\] Divide both sides by ln(6): \[x = \frac{\ln(2)}{\ln(6)}\]

OpenStudy (lukebluefive):

That looks like the answer!

OpenStudy (aaronq):

thanks man! that still doesn't work out when i check it (it equals 2.5). But it's cool, i'll give it another try later today.